For a data set obtained from a sample, n = 100 and x−x- = 44.25. It is known that σ = 5.5.

(a) What is the point estimate of µ?

b) Find α corresponding to a 95% confidence level.

(c) Find 1 − α/2.

(d) Find to 2 decimal places zα/2 for 95% confidence level.
Answer #1

Solution :


Given that,

a

Point estimate = sample mean =     = 44.25

At 95% confidence level the z is ,

b.

= 1 - 95% = 1 - 0.95 = 0.05

c.

/ 2 = 0.05 / 2 = 0.025

d

Z/2 = Z0.025 = 1.96   ( Using z table )